Suppose an investor purchases $10,000 of par value of a TIPS. The real estate (determined at the auction) is 3.8% Assume that at the end of the first six months the CPI-U is 2.4% (annual rate) and at the end of the second six months the CPI-U is 2.8% (annual rate).

i) Compute the inflation-adjusted principal at the end of the first six months.

ii) the coupon payment made to the investor at the end of the first six months

iii) the inflation-adjusted principal at eh end of the second six months

iv) the coupon payment made to the investor at the end of the second six months
Answer #1
  • AT THE END OF FIRST SIX MONTHS
  • Inflation adjustment to the principal at the end of first six months: $10,000 * (0.024 * 6/12) = $120
  • Inflation adjusted principal is 10,000 * 1.012 = $10,120
  • Coupon payment at the end of first six months = one half of the coupon rate (real rate) * inflation adjusted principal at the coupon payment date: 10,120 * (0.038 * 6/12 ) = $192.28
  • AT THE END OF SECOND SIX MONTHS
  • Inflation adjustment to the principal at the end of second six months: $10,120 * (0.028 * 6/12) = $141.68
  • Inflation adjusted principal is 10,120 * 1.014 = $10,261.68
  • Coupon payment at the end of second six months = one half of the coupon rate (real rate) * inflation adjusted principal at the coupon payment date: 10,261.68 * (0.038 * 6/12 ) = $194.97